# A coupled-trajectory approach for decoherence, frustrated hops and internal consistency in surface hopping

### v2

@@ Abstract
In summary, we have identified a new and robust strategy to simulate the electron-nuclear dynamics underlying nonadiabatic processes, which is based on the exact factorization and employs the idea of surface hopping to approximate the nuclear dynamics. The crucial ingredient is the coupling of the trajectories, a concept that emerges naturally in the context of the exact factorization, and is the key to achieve internal consistency, to circumvent overcoherence, to alleviate the frustrated hops, and to produce stable results upon changing the velocity rescaling within surface hopping. The coupling of the trajectories is essential to encode non-local, purely quantum mechanical features in the electronic and nuclear dynamics simulated using a surface hopping scheme, to calculate the quantum momentum and to introduce the energy sharing. Three practical schemes were proposed and tested on full-dimensional photodynamics of fulvene, based on a linear vibronic coupling model. Out of three, two schemes were shown to give robust results for electronic and nuclear properties. All the different flavors of CCT-TSH have been implemented in G-CTMQC [53] and are freely available on GitLab, while the LVC fulvene model has been implemented in QuantumModelLib[54].
Acknowledgements.

### v3

@@ 4 Conclusions
In summary, we have identified a new and robust strategy to simulate the electron-nuclear dynamics underlying nonadiabatic processes, which is based on the exact factorization and employs the idea of surface hopping to approximate the nuclear dynamics. The crucial ingredient is the coupling of the trajectories, a concept that emerges naturally in the context of the exact factorization, and is the key to achieve internal consistency, to circumvent overcoherence, to alleviate the frustrated hops, and to produce stable results upon changing the velocity rescaling within surface hopping.
The coupling of the trajectories is essential to encode non-local, purely quantum mechanical features in the electronic and nuclear dynamics simulated using a surface hopping scheme, to calculate the quantum momentum and to introduce the energy sharing. Indeed, while the energy sharing is an idea grounded in the quantum-mechanical nature of the dynamics that the trajectories represent, its actual implementation remains somehow arbitrary. Three practical schemes were proposed and tested on full-dimensional photodynamics of fulvene and DMABN, based on linear vibronic coupling models. Out of these three flavors of CCT-TSH, two schemes, namely the equity-based and the overlap-based, were shown to give robust results for electronic and nuclear properties while reducing the occurences of frustrated hops, but we expect that alternative formulations might perform equally well.
The deterministic hopping strategy in CCT-TSH, guided directly by the time evolution of the electronic populations obtained from the exact factorization framework, proved to work quite well. By ensuring that nuclear trajectories evolve consistently in the electronic state with the largest population, the method maintains coherence between the electronic and nuclear dynamics. Our results demonstrate that this approach offers a robust and physically motivated alternative to the stochastic hopping schemes, with performance comparable to, or in some cases better than, the fewest-switches variant.
{acknowledgement}
@@
{suppinfo}
We provide as Supporting Information Python script (only the library tkinter is required to run the script) to help generate the input files to run all the calculations reported below, thus, all our data are easily reproducible.